# Autoregressive Semantic Visual Reconstruction Helps VLMs Understand Better

### v1

@@ Abstract
Typical large vision-language models (LVLMs) apply autoregressive supervision solely to textual sequences, without fully incorporating the visual modality into the learning process. This results in three key limitations: (1) an inability to utilize images without accompanying captions, (2) the risk that captions omit critical visual details, and (3) the challenge that certain vision-centric content cannot be adequately conveyed through text. As a result, current LVLMs often prioritize vision-to-language alignment while potentially overlooking fine-grained visual information. While some prior works have explored autoregressive image generation, effectively leveraging autoregressive visual supervision to enhance image understanding remains an open challenge. In this paper, we introduce Autoregressive Semantic Visual Reconstruction (ASVR [^1] ASVR can be pronounced as “as-we-are”), which enables joint learning of visual and textual modalities within a unified autoregressive framework. We show that autoregressively reconstructing the raw visual appearance of images does not enhance and may even impair multimodal understanding. In contrast, autoregressively reconstructing the semantic representation of images consistently improves comprehension. Notably, we find that even when models are given continuous image features as input, they can effectively reconstruct discrete semantic tokens, resulting in stable and consistent improvements across a wide range of multimodal understanding benchmarks. Our approach delivers significant performance gains across varying data scales (556k-2M) and types of LLM bacbones. Specifically, ASVR improves LLaVA-1.5 by 5% in average scores across 14 multimodal benchmarks. The code is available at https://github.com/AlenjandroWang/ASVR.
@@ 1 Introduction
The success of large language models (LLMs) has demonstrated the tremendous potential and scalability of the autoregressive (AR) paradigm. In recent years, extending LLMs’ powerful capabilities to multimodal understanding through bridge-style architectures, exemplified by LLaVA [28, 29, 31], have achieved remarkable performance across vision-language tasks [32, 61, 12, 14, 25, 17, 21]. These models [4, 52, 60, 6, 35, 57], typically adopt a simple yet effective learnable projector to align features from a CLIP-based visual encoder into the text embedding space of LLMs.
However, most of the current large vision-language models (LVLMs) [53, 8, 30, 23] supervise only the textual outputs, overlooking the rich visual modality. Specifically, these models are trained to predict the next token in a text response given both the preceding text and associated images. For example, LLaVA-1.5 [27] represents a single 336$\times$336 image with 576 visual tokens, yet applies no explicit supervision to the visual content. As a result, while these models are multimodal in form, they remain predominantly language-centric in nature, with insufficient attention paid to the visual modality.
To overcome the lack of explicit visual supervision, traditional LVLMs rely on image-caption pairs to associate visual content with language. However, this approach suffers from three critical limitations, as shown in Figure. 1: (1) Although there is a vast amount of image data available online, most images are not accompanied by detailed captions; (2) Even when captions are generated, either manually or by LVLMs, the process is costly, and there remains a risk of omitting critical visual details. The descriptive richness of these captions ultimately constrains the granularity of the model’s visual understanding; (3) Some vision-centric content simply cannot be adequately conveyed through text. As the saying goes, “a picture is worth a thousand words”, the visual modality serves as an independent and expressive channel that captures spatial relationships, textures, complex compositions, and subtle stylistic cues that text alone struggles to express. In summary, the full spectrum of visual detail in an image is difficult to articulate comprehensively through text, and acquiring large-scale, high-quality, fine-grained captions remains both labor-intensive and expensive.
Figure 1: (Left) A simple illustration that reflects the information loss faced by language-centric approaches. (Right) Our proposed Autoregressive Semantic Visual Reconstruction (ASVR) brings significant improvements across various aspects, including General VQA, Visual-centric, Hallucination, and OCR. All the scores are normalized by $x_{\text{norm}}=(x-x_{\text{min}}+10)/(x_{\text{max}}-x_{\text{min}}+10)$.
Recently, several pioneering works have explored unifying visual understanding and generation within the autoregressive paradigm of LLMs [44, 54, 56, 46], where visual tokens are supervised through image generation tasks. However, these studies primarily focus on leveraging visual understanding to enhance generation, rather than investigating the reverse direction. Effectively utilizing autoregressive visual supervision to improve visual understanding remains an open challenge. Most recently, Wang et al. [51] proposed supervising visual outputs via a denoising approach. However, their method relies on external Diffusion Transformer (DiT) modules for visual supervision and lacks a unified framework that aligns visual and textual modalities under a unified supervision scheme.
@@
In this paper, we introduce Autoregressive Semantic Visual Reconstruction (ASVR), a method that enables joint learning of visual and textual modalities within the unified autoregressive framework of LLMs, without relying on any external modules. Specifically, ASVR allows LVLMs to supervise visual outputs by autoregressively predict the next discrete semantic token of input images, which is prepared by a pretrained semantic visual tokenizer [43, 56, 40, 59]. Interestingly, we show that autoregressively reconstructing the raw visual appearance of images does not improve and may even degrade multimodal understanding. In contrast, reconstructing semantic visual representation autoregressively consistently enhances the visual understanding capabilities of LVLMs. Notably, we find that even when models are provided with continuous image features as input, they can effectively reconstruct discrete semantic tokens. This setting even outperforms approaches where both input and output use shared discrete semantic visual tokens, resulting in considerable gains.
@@
Our approach delivers significant and consistent performance gains across varying data scales( LLaVA-1.5-665K [27], LLaVA-Next-779K [30], Bunny-v1_1-data-2M [16]) and model architectures such as Vicuna family [65] as well as Mistral [20]. Specifically, ASVR improves LLaVA-1.5 by 5% in average scores across 14 multimodal benchmarks and the effectiveness is robust across different visual feature types, LLM backbone capacities, data scales, and high-resolution scenarios. These results underscore the importance of explicit semantic visual supervision in training LVLMs. ASVR not only improves visual understanding but also introduces a scalable, unified training strategy, offering a new perspective on autoregressive modeling for multimodal systems.
2 Related Work
@@
The rapid progress in large language models (LLMs)[3, 1, 48, 5, 39, 38] has showcased their strong generalization and remarkable instruction-following capabilities. To further expand these strengths for interpreting and interacting with the world through both visual and linguistic channels. There has been growing interest in Large Vision-Language Models (LVLMs)[28, 27, 30], typically trained using a straightforward two-stage visual instruction tuning paradigm [28], and align visual features extracted by visual encoder with the knowledge and reasoning capabilities of LLMs through the lightweight projector. This process involves jointly training the projector and the LLM on visual instruction datasets, with optional fine-tuning of the visual encoder. However, supervision is limited to text outputs. ASVR introduces a novel autoregressive visual semantic supervision mechanism that encourages the LVLM to reconstruct semantic visual tokens, enhancing its multimodal understanding capabilities.
Visual Autoregression for LVLMs
Some recent approaches [44, 40, 54, 56, 55], introduce autoregressive visual supervision via visual tokenizers, such as VQGAN [11] and VQ-VAE [49], enabling LVLMs to support both multimodal understanding and image generation by predict relevant next visual tokens, which are then decoded into images. In contrast,  ASVR focuses specifically on enhancing the multimodal understanding capability of LVLMs. Rather than generating images,  ASVR employs autoregressive visual supervision to reconstruct semantic visual tokens within the given continuous image features as input. While prior methods are generative,  ASVR adopts the reconstructive approach aimed at promoting perception of visual information.
Reconstructive Objectives for LVLMs
ROSS[50] introduced visual supervision for LVLMs by applying denoising objective to reconstruct visual tokens. In contrast,  ASVR proposes a unified approach by employing autoregressive objective—analogous to that used for text—to reconstruct semantic visual tokens. This design enables seamless integration of visual and textual information under a unified next-token prediction paradigm.
3 Preliminaries
Large Vision Language Models Modeling
To process and represent input sequences from different modalities in a unified manner, Large Vision-Language Models (LVLMs) typically comprise three components: a pre-trained Large Language Model (LLM), a projector commonly implemented as two-layer MLP and a pre-trained visual encoder with semantic aligned.
Given a input RGB image $I\in\mathbb{R}^{H\times W\times 3}$, where $H$ and $W$ denote the image height and width, a pre-trained visual encoder $V_{\xi}$ is first used to extract image features $\mathbf{z}^{I}=V_{\xi}(I)$. These features are then mapped into LLM embedding space via a projector $P_{\phi}$, producing a sequence of visual features: $\mathbf{H}^{I}=P_{\phi}(\mathbf{z}^{I})\in\mathbb{R}^{m\times d}$, where $m=h\times w$ denotes the length of visual features, and $d$ is the embedding dimension of LLM. $\xi$ and $\phi$ are the parameters of the visual encoder and projector, respectively. For a textual input $T\in\mathbb{Z}^{L}$, the LLM’s tokenizer is used to produce a sequence of token indices $\mathbf{x}^{T}=\text{Tokenizer}(T)\in\mathbb{R}^{n}$. These indices are then transformed into textual embeddings via the LLM’s embedding layer $\mathbf{H}^{T}=\text{Embedding}({x}^{T})\in\mathbb{R}^{n\times d}$ where $n$ denotes the sequence length.
The final multimodal inputs are formed by concatenating the visual features and textual embeddings, resulting in $[\mathbf{H}^{I},\mathbf{H}^{T}]\in\mathbb{R}^{(m+n)\times d}$, which is then fed into a causal LLM backbone $L_{\theta}$ with parameters $\theta$ for unified autoregressive modeling:
@@ 3 Preliminaries
Training Framework for LVLMs
LVLM training generally involves two stages [28]: pre-training and instruction tuning. Pre-training aligns different modalities, enabling the model to jointly understand visual and textual inputs. Instruction tuning further enhances generalization across diverse downstream tasks such as Visual Question Answering (VQA).
The training objective is to maximize the the probability of the target textual responses in autoregressive manner, where only textual responses following the $s$-th token position are supervised.
$\mathcal{L}_{\mathrm{AR}}^{\text{text}}(\Theta=\{\theta,\xi,\phi\},T,I)=\frac{-1}{n-s}\sum_{i=s+1}^{n}\log L_{{\theta}}(x_{i}^{T}\mid x_{<i}^{T},\mathbf{H}^{I}),$ (2)
Here, $\Theta$ denotes the parameters of the entire LVLM. During pre-training, only the parameters of the projector $\phi$ are typically updated, while in instruction tuning, the LLM parameters $\theta$ are also finetuned. The visual encoder $v_{\xi}$ may either remain frozen [28, 45] or be jointly optimized [23, 8, 53, 30].
@@ 4 Method
In this section, we introduce ASVR. An overview of the method is provided in Section 4.1, followed by detailed analyses of the visual tokenizer and visual encoder in Sections 4.2 and 4.3, respectively. The training procedure is detailed in Section 4.4. A detailed comparison between the typical LVLMs (LLaVA) and our ASVR is illustrated in Figure 2, highlighting the key innovation of incorporating autoregressive visual supervision to enhance the model’s multimodal understanding capabilities.
@@ 4.1 Overview
Figure 2: Left: the typical LVLM framework exemplified by LLaVA [28]. Right: overview of ASVR’s model architecture and training procedure. The input image and its corresponding text are tokenized into sequences of discrete token indices for unified autoregressive supervision over both visual and textual outputs. For each module, the icon before the slash indicates whether it is frozen or tunable during pre-training, while the icon after the slash indicates its configuration during instruction tuning. "s" and "e" denote the start and end of the text tokens, respectively.
We incorporate autoregressive visual supervision into the typical LVLM’s framework described in Section 3 by extending the next-token prediction paradigm to reconstruct and perceive visual inputs. This unified formulation enables the model to seamlessly integrate visual and textual information—first perceiving, then reasoning—thereby establishing a perceptual foundation for image understanding, alleviating the information loss caused by text-only supervision, and ultimately enhancing the LVLM’s multimodal understanding capabilities.
As illustrated in Figure 2 (b), we employ the visual tokenizer to convert the input image into discrete sequence of visual token indices, serving as visual supervision signals $\mathbf{x}^{I}=\text{Tokenizer\_img}(I)\in\mathbb{R}^{m}$ where $m$ matches the length of the visual features sequence $\mathbf{H}^{I}$ extracted from pre-trained visual encoder and fed into the LLM backbone. The visual head tailored to the visual tokenizer is then trained to predict the next visual token in autoregressive manner, analogous to textual supervision:
$\mathcal{L}_{\mathrm{AR}}^{\text{vision}}(\Theta=\{\theta,\xi,\phi\},I)=\frac{-1}{m}\sum_{i=1}^{m}\log L_{{\theta}}(x_{i}^{I}\mid x_{<i}^{I}),$ (3)
Then our final training objective is combined with $\mathcal{L}_{\mathrm{AR}}^{\text{vision}}$ and $\mathcal{L}_{\mathrm{AR}}^{\text{text}}$, formulated as
$\mathcal{L}_{\mathrm{AR}}(\Theta=\{\theta,\xi,\phi\},I,T)=\mathcal{L}_{\mathrm{AR}}^{\text{vision}}+\mathcal{L}_{\mathrm{AR}}^{\text{text}}$ (4)
This design unifies the learning paradigm across modalities, enabling joint optimization of both vision and language under shared autoregressive objective. Importantly, it also compels the model to first develop coherent visual sensor, which subsequently serves as foundation for more accurate and contextually grounded multimoda understanding.
@@ 4.2 Visual Tokenizer
Visual tokenizer convert input images into one-dimensional sequences of discrete visual codes through vector quantization(VQ) by learning a fixed-size visual codebook, then look up the corresponding features by codes into the codebook as inputs to the LMM. Additionally, the visual tokenizer defines visual supervision targets by determining the granularity and representations of the discrete visual token indices, which play a critical role in the visual reconstruction and perception. There are two type of visual tokenizer.
@@
A visual appearance tokenizer [11, 44] is optimized with the objective of reconstructing the input image, where utilize reconstruction loss typically combining pixel-wise L2 loss [9], LPIPS loss[64] and adversarial loss [19] for reconstruction ability. The resulting sequence of token indices represents a quantized mapping of the image’s pixel-level features. Using Pixel-based tokenizer to provide visual pixel supervision targets will guide the LVLM to focus on low-level pixel feature reconstruction and perception.
@@
A visual semantic tokenizer [40, 56, 59, 43] is is trained to align image features with textual semantics, typically using a contrastive loss [41] to enhance cross-modal alignment. The resulting sequence of token indices represents a quantized mapping of the image’s high-level semantic features. Using Semantic-based tokenizer to provide semantic visual supervision targets will guide the LVLM to focus on semantically meaningful aspects reconstruction and perception of the image, thereby promoting more effective multimodal understanding.

### v2

@@ Abstract
Typical large vision-language models (LVLMs) apply autoregressive supervision primarily to textual responses, without fully exploiting causal learning over rich visual inputs. As a result, these models often emphasize vision-to-language alignment while potentially overlooking fine-grained visual information. While prior work has explored autoregressive image generation, effectively leveraging autoregressive visual supervision to enhance image understanding remains an open challenge. In this paper, we introduce Autoregressive Semantic Visual Reconstruction (ASVR), which enables joint learning of visual and textual modalities within a unified autoregressive framework. ASVR trains models to autoregressively reconstruct the semantic content of input images, which consistently enhances multimodal comprehension. Notably, we show that even when provided with continuous image features as input, models can effectively reconstruct discrete semantic tokens, resulting in stable and consistent improvements across various multimodal understanding benchmarks. ASVR delivers significant performance gains and scalability across varying data scales, visual input, visual supervision and model architectures. In particular, ASVR generally improves baselines by 2-3% across 14 multimodal benchmarks.
Autoregressive Semantic Visual Reconstruction
Helps VLMs Understand Better
Dianyi Wang1,2thanks: Co-first authors.  Wei Song2,3,411footnotemark: 1  Yikun Wang1,2  Siyuan Wang6 Kaicheng Yu3  Zhongyu Wei1,2thanks: Corresponding author.  Jiaqi Wang2,522footnotemark: 2 1Fudan University  2Shanghai Innovation Institute  3AutoLab, Westlake University 4Zhejiang University  5Shanghai AI Lab  6University of Southern California dywang24@m.fudan.edu.cn, songweii@zju.edu.cn
@@ 1 Introduction
The success of large language models (LLMs) has demonstrated the tremendous potential and scalability of the autoregressive (AR) paradigm. Recent advances extending LLMs’ powerful capabilities to multimodal understanding through bridge-style architectures, exemplified by LLaVA (liu2023llava; liu2024llava1d5; liu2024llavanext), have achieved remarkable performance across vision-language tasks (liu2023mmbench; yue2023mmmu; fu2024mme; vqa_v2; POPE; gqa; kembhavi2016diagramworthdozenimages). These models (Qwen-VL; Qwen2VL; yao2024minicpm; chen2024internvl; lu2024deepseekvl; wu2024deepseekvl2), typically adopt a learnable projector to align features from a pretrained visual encoder into the text embedding space of LLMs.
However, most large vision-language models (LVLMs)(wang2024qwen2vlenhancingvisionlanguagemodels; dong2024internlmxcomposer24khdpioneeringlargevisionlanguage; liu2024llava; li2024llavaonevisioneasyvisualtask) supervise only textual outputs through next-token prediction, while overlooking the causal learning of rich visual content within input images. Although this limitation is partially mitigated by training on image–caption pairs that associate visual content with language, the visual modality expresses far more than text alone, capturing spatial relationships, textures, complex compositions, and subtle stylistic cues that language cannot fully convey. For example, LLaVA-1.5liu2023improved represents a single $\times336$ image with 576 visual tokens, which collectively encode substantially more information than the associated caption, yet applies no explicit supervision to this visual content.
Figure 1: (Left) Unified LVLMs aim to implicitly improve visual understanding through additional text-to-image (T2I) objectives. (Mid) ROSS introduces an auxiliary diffusion transformer (DiT) to reconstruct pixel-level image information conditioned on latent representations from the LVLM. (Right) ASVR directly reconstructs the semantic content of the input image, enhancing semantic-level alignment.
While recent unified models have explored integrating visual understanding and generation within the autoregressive paradigm (team2024chameleon; wang2024emu3; wu2024vilau; metamorph), visual tokens are typically supervised only on the output side through visual generation objectives. In contrast, visual tokens fed as inputs are not explicitly supervised to enhance visual understanding. Effectively supervising autoregressive visual inputs to improve fine-grained visual understanding remains an open challenge. Most recently, wang2024ross proposed reconstructing visual inputs via denoising, yet their method relies on external Diffusion Transformer (DiT) modules and lacks a unified next-token prediction framework that explicitly encourage causal learning over detailed visual information and dependencies rather than high-level bidirectional contextual correlations.
@@
In this paper, we introduce Autoregressive Semantic Visual Reconstruction (ASVR), a method that enables joint learning of the visual input and textual output within a unified autoregressive paradigm, without relying on any external modules. Specifically, ASVR allows LVLMs to autoregressively predict the next discrete semantic token of an input image, which is prepared by a pretrained semantic visual tokenizer (song2025dualtoken; wu2024vilau; qu2024tokenflow; xie2024musevl). Autoregressively reconstructing semantic visual representations consistently enhances the visual understanding capabilities and further improve reasoning performance. Notably, we find that models can effectively reconstruct discrete semantic tokens even when provided with continuous image features as input. This setting yields substantial gains over approaches that use shared discrete semantic visual tokens for both input and output, and also outperforms the denoising-based visual reconstruction paradigm (wang2024reconstructivevisualinstructiontuning).
@@
Our approach delivers significant and consistent gains across varying data scales (LLaVA-1.5-665K (liu2023improved), LLaVA-Next-779K (liu2024llava), Bunny-v1_1-data-2M (he2024efficientmultimodallearningdatacentric),LLaVA-OV-4M (li2024llavaonevisioneasyvisualtask)) as well as across model architectures including Vicuna family (zheng2023judgingllmasajudgemtbenchchatbot) and Mistral (jiang2023mistral). Specifically, ASVR improves baselines by 2-3% on average across 14 multimodal benchmarks with improvements remaining robust across visual feature types, LLM backbone capacities, data scales, and high-resolution scenarios. These results underscore the importance of explicit semantic visual supervision in training LVLMs. ASVR not only improves visual understanding but also introduces a scalable, unified training strategy, offering a new perspective on autoregressive modeling for multimodal systems.
2 Preliminaries
@@
To process and represent input from different modalities in a unified manner, LVLMs typically comprise three components: a pre-trained LLM core, a projector commonly implemented as a two-layer MLP, and a pre-trained visual encoder with semantic alignment.
Given a input RGB image $I\in\mathbb{R}^{H\times W\times 3}$, where $H$ and $W$ denote the image height and width, a pre-trained visual encoder $V_{\xi}$ first extracts visual features $\mathbf{z}^{I}=V_{\xi}(I)$. These features are then projected into the LLM embedding space through a projector $P_{\phi}$, yielding a sequence of visual embeddings: $\mathbf{H}^{I}=P_{\phi}(\mathbf{z}^{I})\in\mathbb{R}^{m\times d}$, where $m=h\times w$ denotes the length of visual features, and $d$ is the embedding dimension of LLM. $\xi$ and $\phi$ are the parameters of the visual encoder and projector, respectively. For a textual input $T\in\mathbb{Z}^{L}$, the LLM tokenizer produces a sequence of token indices $\mathbf{x}^{T}=\text{Tokenizer}(T)\in\mathbb{R}^{n}$. which are then mapped into textual embeddings via the LLM’s embedding layer $\mathbf{H}^{T}=\text{Embedding}({x}^{T})\in\mathbb{R}^{n\times d}$ where $n$ denotes the text sequence length.
The multimodal input is formed by concatenating the visual and textual embeddings as $\mathbf{H}^{I},\mathbf{H}^{T}]\in\mathbb{R}^{(m+n)\times d}$, which is then fed into a causal LLM backbone $L_{\theta}$ with parameters $\theta$ for unified autoregressive modeling:
@@
Training Pipeline for LVLMs
LVLMs training typically follows a two-stage paradigm (liu2023llava): pre-training and instruction tuning. During pre-training, the model learns to align visual and textual modalities, enabling joint understanding of multimodal inputs. Instruction tuning further enhances generalization across diverse downstream tasks such as visual question answering (VQA). The training objective is to maximize the likelihood of target textual responses in an autoregressive manner, with supervision applied only to textual responses. In practice, pre-training usually updates only the projector parameters $\phi$, while instruction tuning additionally fine-tunes the LLM parameters $\theta$. The visual encoder $v_{\xi}$ may either remain frozen (liu2023llava; tong2024cambrian1fullyopenvisioncentric) or be jointly optimized (li2024llavaonevisioneasyvisualtask; dong2024internlmxcomposer24khdpioneeringlargevisionlanguage; wang2024qwen2vlenhancingvisionlanguagemodels; liu2024llava).
@@ 3 Method
In this section, we introduce ASVR which learns autoregressive modeling of textual responses while simultaneously reconstructs visual inputs autoregressively to enhance visual understanding. An overview of the method is provided in Section 3.1, followed by detailed descriptions of the visual tokenizer and visual encoder in Sections 3.2 and 3.3, respectively. The training procedure is described in Section 3.4. Figure 2 provides a detailed comparison between conventional LVLMs (e.g., LLaVA) and our ASVR, highlighting the key innovation of incorporating autoregressive visual input supervision to enhance multimodal understanding.
@@ 3.1 Overview
Figure 2: Left: The typical LVLM paradigm such as LLaVA (liu2023llava). Right: Overview of ASVR’s architecture and training pipeline. The input image and text are tokenized into sequences of discrete token indices enabling unified autoregressive supervision over both visual inputs and textual outputs. For each module, the icon before the slash indicates whether it is frozen or tunable during pre-training, while the icon after the slash indicates its configuration during instruction tuning. "s" and "e" denote the start and end of the text tokens, respectively.
We incorporate autoregressive visual supervision into typical LVLMs described in Section 2 by extending the next-token prediction paradigm to reconstruct visual inputs. This unified formulation enables the model to seamlessly integrate visual and textual information, thereby establishing a perceptual foundation for image understanding, alleviating information loss caused by text-only supervision, and ultimately enhancing multimodal understanding and reasoning capabilities.
As illustrated in Figure 2(b), we employ a visual tokenizer to convert the input image $I$ into a discrete sequence of visual token indices $\mathbf{x}^{I}\in\mathbb{R}^{m}$, which serve as visual supervision signals and $m$ matches the length of the visual features sequence $\mathbf{H}^{I}$ extracted by the pre-trained visual encoder and fed into the LLM backbone. A visual prediction head tailored to the visual tokenizer is trained to autoregressively predict the next visual token, analogous to textual supervision:
$\mathcal{L}_{\mathrm{AR}}^{\text{vision}}(\Theta,I)=-\frac{1}{m}\sum_{i=1}^{m}\log L_{{\theta}}(x_{i}^{I}\mid x_{<i}^{I}),$ (2)
where $\Theta=\{\theta,\xi,\phi\}$ denotes the parameters of the LLM backbone, visual encoder and projector. The final training objective jointly optimizes visual and textual autoregressive losses, $\mathcal{L}_{\mathrm{AR}}^{\text{vision}}$ and $\mathcal{L}_{\mathrm{AR}}^{\text{text}}$:
$\mathcal{L}_{\mathrm{AR}}(\Theta,I,T)=\mathcal{L}_{\mathrm{AR}}^{\text{vision}}+\mathcal{L}_{\mathrm{AR}}^{\text{text}}$ (3)
This design unifies the learning paradigm across modalities, and encourages the model to first develop a coherent visual understanding, which subsequently serves as a foundation for more accurate and contextually grounded multimoda reasoning.
@@ 3.2 Visual Tokenizer
Unlike continuous visual encoders, a visual tokenizer converts an input image into a one-dimensional sequence of discrete visual codes through vector quantization (VQ) by learning a fixed-size visual codebook. The corresponding embeddings are retrieved from the codebook based on these codes and used as inputs to the LLM backbone. In this work, we utilize such a visual tokenizer to obtain discrete visual token representations as supervision targets for visual reconstruction. Existing visual tokenizers can be broadly categorized into two types.
@@
A visual appearance tokenizer (esser2021vqgan; team2024chameleon) is optimized with the objective of reconstructing the appearance of the input image, typically using a combination of pixel-wise L2 loss (dosovitskiy2016L2), LPIPS losszhang2018lpips and adversarial loss isola2017gan. The resulting sequence of token indices corresponds to a quantized mapping of the image’s pixel-level features. Using appearance tokenizers to provide pixel-level supervision will encourage LVLMs to focus on low-level visual feature reconstruction.
@@
A visual semantic tokenizer (qu2024tokenflow; wu2024vilau; xie2024musevl; song2025dualtoken) is trained to align image features with textual semantics, typically using a contrastive loss (radford2021clip) to enhance cross-modal alignment. The resulting sequence of token indices represents a quantized mapping of the image’s high-level semantic features. Using semantic tokenizers to provide semantic visual supervision will guide LVLMs to focus on semantically meaningful structures and relationships of the image, thereby promoting more effective multimodal understanding.